OpenStudy (anonymous):

Solve. x2 + 15x = –56 a. x = 8 and x = 7 b. x = –8 and x = –7 c. x = –8 and x = 7 d. x = 8 and x = –7 I think its b, am i right?

OpenStudy (anonymous):

x^2+15x+56 (x+8)(x+7) x+8=0 -8 = -8 x=-8 x+7=0 -7 -7 x=-7 so x=-8 and x=-7 so yes its b

OpenStudy (anonymous):

thank u